OpenStudy (anonymous):

In how many ways can 10 people sit on a park bench that only seats 3? Order is important.

OpenStudy (raden):

it means : 10P3 = 10!/(10-3)! = 10!/7! = (10 * 9 * 8 * 7!)/7! cancel out the 7! giving us = 10 * 9 * 8 = 720 ways